OpenStudy (anonymous):

The vertices of the trapezoid are the origin, A(4m, 4n), B(4q, 4n), and C(4p, 0). Find the midpoint of the midsegment of the trapezoid.

OpenStudy (anonymous):

a. (2q, 2n) c. (m + q + p, n) b. (m + q + p, 2n) d. (2m + 2p, 2n)

OpenStudy (mathstudent55):

Have you tried graphing it?

OpenStudy (anonymous):

It came with a graph but this confuses me all to heck

OpenStudy (mathstudent55):

No, just the opposite. The graph will help because it will show you which sides you need to get the midpoints of.

OpenStudy (mathstudent55):

Do you know what the midsegment of a trapezoid is?

OpenStudy (thomas5267):

What exactly is the midpoint of the midsegment?

OpenStudy (anonymous):

I know what it is but it's more of the coordinates just being confusing

OpenStudy (mathstudent55):

The midsegment of a trapezoid is the segment that connects the midpoints of the non-parallel sides of the trapezoid.

OpenStudy (mathstudent55):

By looking at a graph, you can easily tell which are the two non-parallel sides.

OpenStudy (mathstudent55):

Then find the midpoint of each of the two non-parallel sides.

OpenStudy (mathstudent55):

Do you know how to find the midpoint of a segment given its endpoints?

OpenStudy (anonymous):

Sadly I don't

OpenStudy (mathstudent55):

That's not a hard thing to do. Take the x-coordinates and find their average. Take the y-coordinates and find their average.

OpenStudy (mathstudent55):

For example, the midpoint of segement AB with coordinates A(0, 6) and B(8, 12) is \(\left(\dfrac{0 + 8}{2}, \dfrac{6 + 12}{2} \right) = (4, 9)\)

OpenStudy (mathstudent55):

Does your trapezoid look somewhat like this one?